OpenStudy (anonymous):

An airplane travels 2100 km at a speed of 800 km/hr, and then encounters a tailwind that boosts its speed to 1000 km/hr for the next 1800 km. What was the average speed of the plane for this trip?

OpenStudy (shane_b):

\[Speed_{avg}=\frac{distance}{time}\]You know the total distance will be:\[2100km+1800km=3900km\] Now you need to find the total time. First leg:\[\frac{2100km}{800km/h}=2.625h\]Second leg:\[\frac{1800km}{1000km/h}=1.8h\]Total time:\[t=1.8h+2.625h=4.425h\] Now find the average speed:\[Speed_{avg}=\frac{3900km}{4.25h}=881.36km/h\]